OpenStudy (anonymous):

let f(x)=sqrt(x(x-1)) whats the domain of f

OpenStudy (anonymous):

We can't take the square root of a negative number, so we find the inequality: x^2-x>=0 X^2>=x which is true for x<=0 and for x>=1. And that is your domain.

OpenStudy (anonymous):

so how do you write it???

OpenStudy (anonymous):

D=????

OpenStudy (anonymous):

there are several notations you can use, this is one: {x|x<=0 OR x>=1} or x<=0 V x>=1 is another.

OpenStudy (anonymous):

@Thomas9 "We can't take the square root of a negative number" Actually you can, it's just that the sqrt sign as written refers to the non-negative square root. Good answer though.

OpenStudy (anonymous):

i know, this question wouldn't be any fun though if you consider complex numbers right?

OpenStudy (anonymous):

x^2-x must be non negative. so domain is(-\[\infty\],0]and[1,\[\infty\])